OpenStudy (anonymous):

Which of the following is a rational number? square root 97square root 98, square root 99, square root 100 square root 97 square root 98 square root 99 square root 100

OpenStudy (usukidoll):

a rational number can be expressed as a fraction an irritation number can't be expressed as a fraction

OpenStudy (usukidoll):

*irrational

OpenStudy (usukidoll):

examples of rational numbers \[\LARGE \frac{1}{2}, \frac{3}{1}\] Examples of irrational numbers \[\LARGE \pi, e, \sqrt{3}\]

OpenStudy (usukidoll):

let's start with the last one. what is the square root of 100?

OpenStudy (anonymous):

10 @UsukiDoll

OpenStudy (usukidoll):

yay so that's a rational number. I can express that as a fraction which is \[\large \frac{10}{1} \]

OpenStudy (anonymous):

right @UsukiDoll

OpenStudy (usukidoll):

I don't think there are any more choices for rational numbers... 97 is a prime number and that's in a square root . that's irrational

OpenStudy (usukidoll):

and even if we did simplify the square root of 98 and 99, we will still have a square root left order...

OpenStudy (anonymous):

so 97s the awnser @UsukiDoll

OpenStudy (usukidoll):

no for 2 reasons 97 is a prime number 97 is inside the square root and can't be expressed as a fraction .

OpenStudy (usukidoll):

but not all square roots are irrational ... so we need to be careful on this. if we are left with a square root that can be simplified further .. it becomes rational

OpenStudy (usukidoll):

for example \[\LARGE \sqrt{2}\] is irrational but \[\LARGE \sqrt{4}=2\] is rational

OpenStudy (usukidoll):

there could be some more rational numbers lurking

OpenStudy (usukidoll):

or not...

OpenStudy (usukidoll):

 take the square root of 49 and square root of 9